OpenStudy (anonymous):

If (9^2)p = 9^8, what is the value of p? 2 3 4 6

OpenStudy (danjs):

is p also a power of that 9^2 term?

OpenStudy (anonymous):

Yes @DanJS

OpenStudy (danjs):

\[[9^2]^p = 9^8\]

OpenStudy (danjs):

recall, a power raised to a power is simplified by multiplying the powers [9^2]^p = 9^(2*p)

OpenStudy (danjs):

Since you have the same bases raised to different powers, you can equate the powers from both sides 2p = 8

OpenStudy (danjs):

Or if you need work, take the log base 9 of both sides, move the powers down in front...

OpenStudy (anonymous):

So its D? Or did I do my math wrong?

OpenStudy (danjs):

2p = 8

OpenStudy (danjs):

\[9^{2p}=9^8\] \[\log_{9}9^{2p} = \log_{9}9^8 \] \[(2p)*\log_{9}9 = 8 * \log_{9}9\] \[2p * 1 = 8*1~~~~~~~~~~~;p=4\]